OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y-4=M(x-(-3))?

OpenStudy (jdoe0001):

yes.... .m = slope just get the slope and "stick it" in there :)

OpenStudy (jdoe0001):

well.... hmm when using the coordinates you'd want to use the coordinates from the same point

OpenStudy (anonymous):

1-2/2-(-3)

OpenStudy (anonymous):

-1/5?

OpenStudy (jdoe0001):

so if you use -3, you'd need to use 2 if you use 2, you'd need to use 1 or if you use the "x" of the 1st point, you'd have to use the "y" for the same 1st one

OpenStudy (jdoe0001):

yeap -1/5 = m= slope \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{1}}-{\color{blue}{ 2}}}{{\color{red}{ 2}}-{\color{red}{ (-3)}}}\implies \cfrac{-1}{2+3}\implies -\cfrac{1}{5} \\ \quad \\ y-{\color{blue}{ 2}}={\color{green}{ -\cfrac{1}{5}}}(x-{\color{red}{ (-3)}})\qquad or \\ \quad \\ y-{\color{blue}{ 1}}={\color{green}{ -\cfrac{1}{5}}}(x-{\color{red}{2}})\)

OpenStudy (anonymous):

y-2=-1/5(x-(-3))

OpenStudy (anonymous):

or y-2=-5(x+3)

OpenStudy (jdoe0001):

yeap well \(\bf y-{\color{blue}{ 2}}={\color{green}{ -\cfrac{1}{5}}}(x-{\color{red}{ (-3)}})\implies y-{\color{blue}{ 2}}={\color{green}{ -\cfrac{1}{5}}}(x+{\color{red}{ 3}})\)

OpenStudy (anonymous):

-1/5=-5

OpenStudy (anonymous):

simplifies it

OpenStudy (anonymous):

s###

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

dur

OpenStudy (jdoe0001):

hmm \(\bf -\cfrac{5}{1}\iff -5\qquad \qquad -\cfrac{1}{5}\iff -\cfrac{1}{5}\)

OpenStudy (anonymous):

welp, Thanks alot man

OpenStudy (anonymous):

g2g

OpenStudy (jdoe0001):

yw